4y-x=5+2y 3x+7y=24 use
Answer by LinnW(1048)

4y-x=5+2y 3x+7y=24 use
rewrite 4y-x=5+2y
subtract 2y from each side
2y-x=5
-x+2y=5
_
So we now have
-x+2y=5
3x+7y=24
multiply the top equation by 3
-3x+6y=15
3x+7y=24
add the equations
13y=39
divide each side by 13
y = 3
substitute 3 for y in -x+2y=5
-x+2(3)=5
-x+6=5
add -6 to each side
-x = -1
x = 1
substitute 1 for x and 3 for y in the original equations
4y-x=5+2y
4(3)-1?=5+2(3)
12-1?=5+6
11=11
_
3x+7y=24
3(1)+7(3)?=24
3+21?=24
24=24